# An Overview and Discussion on Using Large Language Models for Implementation Generation of Solutions to Open-Ended Problems

### v1

@@ III-C Knowledge Retrieval
f) Self-Critique Methods/Evaluation- and Verification-Focused Methods: To improve reliability and reduce factual inaccuracies, Chain-of-Verification (CoVe) [103] uses a four-step process: (1) generating an initial response, (2) formulating verification questions to identify potential errors or inconsistencies, (3) answering these questions to produce supporting evidence or rationale, and (4) revising the original response based on validated findings. CoVe has demonstrated over 10% performance improvements compared to basic prompting and Chain-of-Thought (CoT) methods in both context-free and contextual question-answering tasks.
@@
In summary, self-reflection techniques, e.g., Implicit RAG and MP, emphasize reflective reasoning for deepening understanding and clarity before producing an answer, while self-critique methods, i.e. CoVe and VE, concentrate on verifying and refining initial outputs to reduce inaccuracies. Implicit RAG and MP differ in execution: Implicit RAG systematically retrieves the most relevant textual evidence for enhanced context, whereas MP focuses on iterative introspection and confidence evaluation. CoVe and VE diverge in methodology, CoVe generates verification queries for self-checking, whereas VE specifically pinpoints uncertain outputs and edits them using external knowledge.
@@ III-D Reinforcement Learning
Reinforcement Learning (RL) provides a systematic framework for refining LLM behavior by guiding models toward desired objectives through iterative feedback and carefully designed reward signals [105]. There are six main components: agent, environment, state, action, reward, and policy [106]. To apply RL for fine-tuning LLMs, the first step maps the six components to the LLM framework: the LLM represents the policy, the current textual sequence is the state, and based on this state, the LLM generates an action, the next token. This action updates the state, creating a new state that incorporates the newly added token. After generating a complete textual sequence, a reward is determined by assessing the quality of the LLM output. This reward can be used to train a pre-trained reward model or can be directly integrated into the alignment process to guide the behavior of the model.
@@
[108] pioneered fine-tuning models like InstructGPT using human feedback to better adhere to user instructions. Building on this approach, [109] and [110] explored reward modeling and methods to address challenges such as length bias, ensuring outputs are concise and aligned with human expectations. Frameworks like trlX [111] and high-quality datasets introduced by [112] have scaled RLHF applications, improving the performance of large language models (LLMs) in tasks such as summarization, translation, and dialogue generation. Summarization tasks, for example, leverage reinforcement learning (RL) through both extractive and abstractive methods; extractive summarization selects key sentences from the source, while abstractive summarization generates novel sentences to convey the essence of the content [113]. RL optimizes summarization by using rewards based on metrics like ROUGE to iteratively enhance the quality of outputs. Policy optimization, on the other hand, employs pairwise feedback, comparing response pairs to align LLM outputs with human preferences. Techniques such as Pairwise Proximal Policy Optimization simplify the process by directly operating on comparative rewards, avoiding complexities like value function estimation and normalization [114].
PPO algorithms iteratively adjust the weights of a model to maximize the expected reward [107]. Central to this process is the collection of human feedback, which is critical in training reward models. Studies, such as Skywork-Reward [115] and TÜLU-V2-mix [116], utilize human preferences by curating datasets of ranked examples, enabling models to align more effectively with human judgments. [117] introduces tool-augmented reward modeling, integrating external resources like calculators and search engines to refine alignment. Recent generative reward models use synthetic preferences, which are artificially created by sampling and ranking model outputs using a base preference model, to reduce reliance on extensive human feedback. [118] examined efficient methods for collecting pairwise human preferences, optimizing reward model design within RLHF frameworks. Additionally, research on over-optimization risks underscores the importance of balanced training to prevent performance degradation [119]. [114] propose novel pairwise feedback pipelines that improve preference learning and policy optimization by comparing response pairs to better capture human preferences.
@@
RL from AI Feedback (RLAIF) is a training method designed to replace human evaluators with AI systems, offering better scalability and consistency by mitigating the variability of human judgment [121]. In RLAIF, a Reward Model (RM) is trained using preference labels generated by a Large Language Model (LLM). These labels are transformed into a probability distribution through a softmax function and optimized via cross-entropy loss, enabling the RM to guide the training of the target AI model [122]. Various approaches have been proposed to address the specific challenges of RLAIF. One strategy distills AI feedback to train reward models, leveraging AI-generated insights to fine-tune reward systems and create scalable feedback mechanisms. For example, UltraFeedback compiles a large-scale dataset of over one million GPT-4 feedback annotations on 250,000 user-assistant conversations to train reward models [112]. Magpie employs a self-synthesis method, where an aligned LLM generates large-scale alignment data that fine-tunes reward models [123]. HelpSteer2 introduces a permissively licensed preference dataset to train reward models, demonstrating improved alignment with human preferences [124]. Another approach focuses on prompting LLMs to function as reward functions, directly guiding model training through reward scores, as seen in Exploring with LLMs (ELLM) Rewards [125]. Additional work, such as Reward Design with Language Models, emphasizes constructing reward mechanisms that align model outputs with desired outcomes by leveraging LLM capabilities [126]. Self-supervised feedback mechanisms have also been explored; for instance, the Eureka framework introduces a novel approach to reward optimization through self-generated feedback loops [127]. Self-rewarding systems, including Self-Refined LLMs  [128] and Self-Rewarding Language Models (SRLM)  [129], enable iterative refinement of model outputs based on their own evaluations.
@@
RLAIF remains less widely adopted compared to RLHF. This discrepancy stems from challenges, such as difficulties in achieving alignment and the risk of propagating biases inherent in AI-generated feedback [112, 127]. These challenges can create feedback loops that amplify existing biases, constraining model diversity and limiting its ability to generalize effectively [129]. Moreover, the absence of human evaluators in RLAIF can result in a lack of nuance, leading to a narrower latent space influenced by the biases of the training AI [128].
@@
Exploration techniques in RL involves seeking new information to improve future decisions, whereas exploitation capitalizes on current knowledge to maximize immediate rewards [130]. In these algorithms, each action decision can be made stochastic via epsilon-greedy  [131] or entropy regularization  [132] to ensure diverse coverage of the environment, but excessive exploration can be inefficient. Traditional approaches, such as epsilon-greedy  [133] and Boltzmann exploration  [134], introduce randomness without leveraging prior knowledge, slowing convergence. Recent methods, like, ExploRLLM [135] uses LLMs hierarchically to generate high-level plans and low-level affordance-based policies to efficiently explore high-value states while minimizing reliance on step-by-step LLM invocation. This structured approach enhances efficiency but struggles with adaptability in open-ended domains [136].
@@
Soft RLLF integrates natural language as logical feedback to balance exploration and exploitation, enabling improved performance in reasoning tasks such as negation understanding and logical consistency in high-stakes applications [137]. This is achieved by encoding logical consistency checks and negation handling into the learning process, utilizing feedback loops to iteratively refine the agent’s decision-making. However, its effectiveness diminishes when tackling problems requiring broader adaptability and creativity, as it is optimized for structured reasoning [137]. Another recent approach, LLM+Exp [138], employs a dual-LLM framework: one LLM analyzes action-reward trajectories to derive exploration strategies, while the other dynamically adjusts action probabilities to refine future decisions. Action-reward trajectories represent sequences of actions taken by an agent and the corresponding rewards, offering insights into the learning process. Action probabilities define the likelihood of selecting specific actions based on learned patterns and anticipated outcomes. While this adaptive approach excels in structured environments, it faces scalability issues and struggles to generalize effectively to unpredictable or unstructured tasks. Guided Pretraining RL [125] incorporates LLMs to influence exploration, providing contextual background knowledge to prioritize relevant actions and improve sample efficiency. The method combines LLM-generated contextually relevant trajectories with reinforcement learning, allowing the agent to pretrain on meaningful, structured action sequences before fine-tuning its policies in the target environment. However, it falls short in addressing the complexity and variability of problems that demand broader generalization and creative reasoning.
@@
Direct Preference Optimization (DPO) addresses the limitations of RLHL/PPO, which necessitates meticulous oversight and significant computational resources due to the initial phase to train a reward model using a preference dataset, followed by training an RL policy with the pre-trained reward model serving as the environment. DPO offers a simpler alternative by directly optimizing LLM parameters using preference data, bypassing the need for a reward model [139]. DPO relies on a preference loss function trained on datasets of paired human preferences (e.g., “Response A is better than Response B”). Several extensions to DPO improve upon this baseline. For instance, DPOP [140] (also termed DPO-positive) introduces a margin-based term to prevent rewarding both preferred and disfavored outputs concurrently, thereby improving performance on tasks with small edit distances. Specifically, the margin-based term in DPOP introduces a penalty for assigning high probabilities to both preferred and disfavored outputs, ensuring that the model distinctly favors the preferred response to improve task performance. Iterative DPO [129] (also known as online DPO) mitigates distribution shifts by continually updating the policy on newly generated responses, an advantage over vanilla DPO, which can overfit to a narrower distribution. Meanwhile, $\beta$-DPO [141] adaptively tunes the regularization term based on the data quality, making it more robust to noisy preferences. Stepwise DPO (sDPO) [142] partitions the preference dataset to perform incremental updates, leveraging a stronger intermediate reference model at each phase.
DPO methods are advantageous for structured problem-solving, like in creative writing or complex reasoning because they can directly incorporate human preferences and avoid undesired behavior without heavily relying on large-scale reward modeling or complex RL training loops [139]. However, a recurring drawback is their sensitivity to distribution shifts, e.g., when the model starts generating out-of-domain responses, alignment performance can drop unless the reference model or preference data is iteratively updated [143]. Moreover, purely relying on pairwise or setwise human judgments can still introduce label noise or ambiguity, especially for creative or unstructured tasks [144]. Despite these limitations, DPO-based techniques are promising for balancing helpfulness and correctness in open-ended LLM outputs  [145].
@@
Actor-critic methods, such as Advantage Actor-Critic (A2C) and Deep Deterministic Policy Gradient (DDPG), have been effectively adapted to optimize prompts for large language models (LLMs). Frameworks like Prompt Actor-Critic Editing (PACE) [151] employ an iterative process where the actor (the LLM) generates a response $a$ based on a prompt $p$ and input $X$. This process is formalized as
@@
    Adaptive process: It includes the sequence of automated activities performed to create an implementation. It requires devising new means to predict the expected outcomes of the available activities, selecting and adapting an activity to the current context, understanding the degree to which the sequence advances towards creating an implementation, and learning new knowledge available during the process. Also, when addressing collaboration between humans and LLMs to tackle unexpected challenges, such as handling zero-day attacks, the process necessitates reasoning, understanding of prior instructions, and intuitive decision-making within the context of new parameters and constraints. To automate this process, exploring reasoning techniques, including deductive reasoning, inductive reasoning, analogical reasoning, common sense reasoning, tree-of-thoughts, multiple chains of thought, causal reasoning, heuristic reasoning, and symbolic reasoning, is required. Among these, the primary human thought process often involves mapping the current problem to a previously encountered one or identifying similarities with analogous problems, like in analogical reasoning. Consequently, an effective approach to problem modeling could involve neuro-symbolic representations that allow LLMs to dynamically learn and adapt in real-time. Techniques such as grokking, which enable models to discover relationships and patterns through iterative refinement, and masked LLMs are promising methods to achieve this goal. These approaches empower the model to derive connections on the fly, effectively merging learned representations with reasoning capabilities.

### v2

@@ III-C Retrieval-Augmented Generation
f) Self-Critique Methods/Evaluation- and Verification-Focused Methods: To enhance reliability and reduce factual inaccuracies in automated reasoning, self-critique methods have emerged as critical tools. These methods address challenges in producing consistent, accurate outputs by systematically verifying and refining initial responses. Chain-of-Verification (CoVe) [103] uses a four-step process: (1) generating an initial response, (2) formulating verification questions to identify potential errors or inconsistencies, (3) answering these questions to produce supporting evidence or rationale, and (4) revising the original response based on validated findings. CoVe has demonstrated over 10% performance improvements compared to basic prompting and Chain-of-Thought (CoT) methods in both context-free and contextual question-answering tasks.
@@
In summary, self-critique methods, i.e., CoVe and VE, concentrate on verifying and refining initial outputs to reduce inaccuracies, while self-reflection techniques, e.g., Implicit RAG and MP, emphasize reflective reasoning for deepening understanding and clarity before producing an answer. CoVe and VE diverge in methodology: CoVe generates verification queries for self-checking, whereas VE specifically pinpoints uncertain outputs and edits them using external knowledge.
@@ III-D Reinforcement Learning
Reinforcement Learning (RL) provides a systematic framework for refining LLM behavior by guiding models toward desired objectives through iterative feedback and carefully designed reward signals from human feedback, automated metrics, or a pre-trained reward model [105]. There are six main components: agent, environment, state, action, reward, and policy [106]. To apply RL for fine-tuning LLMs, the first step maps the six components to the LLM framework: the LLM represents the policy, while the current textual sequence is the state, and based on this state, the LLM generates an action, the next token. This action updates the state, creating a new state that incorporates the newly added token. After generating a complete textual sequence, a reward is determined by assessing the quality of the LLM output. This reward can be used to train a pre-trained reward model or can be directly integrated into the alignment process to guide the behavior of the model.
@@
[108] created fine-tuned models like InstructGPT using human feedback to better adhere to user instructions. Similarly, [109] and [110] explored reward modeling and methods to address challenges such as length bias, ensuring outputs are concise and aligned with human expectations. Frameworks like trlX [111] and high-quality datasets introduced by [112] have scaled RLHF applications, improving the performance of LLMs in tasks such as summarization, translation, and dialogue generation. Summarization tasks, for example, leverage reinforcement learning (RL) through both extractive and abstractive methods; extractive summarization selects key sentences from the source, while abstractive summarization generates novel sentences to convey the essence of the content [113]. RL optimizes summarization by using rewards based on metrics like ROUGE to iteratively enhance the quality of outputs. Policy optimization, on the other hand, employs pairwise feedback, comparing response pairs to align LLM outputs with human preferences. Techniques such as Pairwise Proximal Policy Optimization simplify the process by directly operating on comparative rewards, avoiding complexities like value function estimation and normalization [114].
PPO algorithms iteratively adjust the weights of a model to maximize the expected reward [107]. Central to this process is the collection of human feedback, which is critical in training reward models. Studies, such as Skywork-Reward [115] and TÜLU-V2-mix [116], utilize human preferences by curating datasets of ranked examples, enabling models to align more effectively with human judgments. Additionally,  [117] introduces tool-augmented reward modeling, integrating external resources like calculators and search engines to refine alignment. Recent generative reward models use synthetic preferences, which are artificially created by sampling and ranking model outputs using a base preference model, to reduce reliance on extensive human feedback.  [118] examined efficient methods for collecting pairwise human preferences, optimizing reward model design within RLHF frameworks. Additionally, research on over-optimization risks underscores the importance of balanced training to prevent performance degradation [119]. [114] propose novel pairwise feedback pipelines that improve preference learning and policy optimization by comparing response pairs to better capture human preferences.
@@
RL from AI Feedback (RLAIF) is a training method designed to replace human evaluators with AI systems, offering better scalability and consistency by mitigating the variability of human judgment [121]. In RLAIF, a Reward Model (RM) is trained using preference labels generated by an LLM. These labels are transformed into a probability distribution through a softmax function and optimized via cross-entropy loss, enabling the RM to guide the training of the target AI model [122]. Various approaches have been proposed to address the specific challenges of RLAIF. For example, UltraFeedback compiles a large-scale dataset of over one million GPT-4 feedback annotations on 250,000 user-assistant conversations to train reward models [112]. Similarly, Magpie employs a self-synthesis method, where an aligned LLM generates large-scale alignment data that fine-tunes reward models [123]. HelpSteer2 introduces a permissively licensed preference dataset to train reward models, demonstrating improved alignment with human preferences [124]. Another approach focuses on prompting LLMs to function as reward functions, directly guiding model training through reward scores, as seen in Exploring with LLMs (ELLM) Rewards [125]. Additional work, such as Reward Design with Language Models, emphasizes constructing reward mechanisms that align model outputs with desired outcomes by leveraging LLM capabilities [126]. Self-supervised feedback mechanisms have also been explored; for instance, the Eureka framework introduces a novel approach to reward optimization through self-generated feedback loops [127]. Self-rewarding systems, including Self-Refined LLMs  [128] and Self-Rewarding Language Models (SRLM)  [129], enable iterative refinement of model outputs based on their own evaluations.
@@
Despite its potential, RLAIF remains less widely adopted compared to RLHF. This discrepancy stems from challenges, such as difficulties in achieving alignment and the risk of propagating biases inherent in AI-generated feedback [112, 127]. These challenges can create feedback loops that amplify existing biases, constraining model diversity and limiting its ability to generalize effectively [129]. Moreover, the absence of human evaluators in RLAIF can result in a lack of nuance, leading to a narrower latent space influenced by the biases of the training AI [128].
@@
Exploration techniques in RL involves seeking new information to improve future decisions, whereas exploitation capitalizes on current knowledge to maximize immediate rewards [130]. In these algorithms, each action decision can be made stochastic via epsilon-greedy  [131] or entropy regularization  [132] to ensure diverse coverage of the environment, but excessive exploration can be inefficient. Traditional approaches, such as epsilon-greedy  [133] and Boltzmann exploration  [134], introduce randomness without leveraging prior knowledge, slowing convergence. Recent methods, like, ExploRLLM [135] presents a hierarchical reinforcement learning framework that combines the strengths of LLMs and affordance-based policies. In this approach, LLMs generate high-level plans to outline strategic goals, while affordance-based policies, which identify actionable possibilities within the environment, execute specific actions to achieve those goals. This method improves exploration efficiency by prioritizing high-value states and reducing reliance on frequent LLM invocations. Despite its effectiveness in structured environments, the approach faces challenges in adapting to dynamic and open-ended domains [136].
@@
Soft RLLF [137] integrates natural language as logical feedback to balance exploration and exploitation, enabling improved performance in reasoning tasks such as negation understanding and logical consistency in high-stakes applications. This is achieved by encoding logical consistency checks and negation handling into the learning process, utilizing feedback loops to iteratively refine the agent’s decision-making. However, its effectiveness diminishes when tackling problems requiring broader adaptability and creativity, as it is optimized for structured reasoning [137]. Another recent approach, LLM+Exp [138], employs a dual-LLM framework: one LLM analyzes action-reward trajectories to derive exploration strategies, while the other dynamically adjusts action probabilities to refine future decisions. Action-reward trajectories represent sequences of actions taken by an agent and the corresponding rewards, offering insights into the learning process. Action probabilities define the likelihood of selecting specific actions based on learned patterns and anticipated outcomes. While this adaptive approach excels in structured environments, it faces scalability issues and struggles to generalize effectively to unpredictable or unstructured tasks. Guided Pretraining RL [125] leverages LLMs to enhance exploration by providing contextual background knowledge. This knowledge helps prioritize relevant actions, improving sample efficiency—allowing the agent to learn more effectively from fewer interactions. The method involves generating structured trajectories, which represent meaningful sequences of actions related to the task, using LLMs. These trajectories are used to pretrain the agent, providing a solid foundation before fine-tuning its policies in the target environment.
While this approach excels in providing structure and reducing exploration costs, it struggles with tasks that require broader adaptability and creative problem-solving. The reliance on predefined trajectories limits its ability to generalize to highly variable or unpredictable environments, where more flexible reasoning is needed.
@@
Direct Preference Optimization (DPO) addresses the limitations of RLHF/PPO, which necessitates meticulous oversight and significant computational resources due to the initial phase to train a reward model using a preference dataset, followed by training an RL policy with the pre-trained reward model serving as the environment. DPO offers a simpler alternative by directly optimizing LLM parameters using preference data, bypassing the need for a reward model [139]. DPO relies on a preference loss function trained on datasets of paired human preferences (e.g., ”Response A is better than Response B”). Several extensions to DPO improve upon this baseline. For instance, DPOP [140] (also termed DPO-positive) introduces a margin-based term to prevent rewarding both preferred and disfavored outputs concurrently, thereby improving performance on tasks with small edit distances. Specifically, the margin-based term in DPOP introduces a penalty for assigning high probabilities to both preferred and disfavored outputs, ensuring that the model distinctly favors the preferred response to improve task performance. Iterative DPO [129] (also known as online DPO) mitigates distribution shifts by continually updating the policy on newly generated responses, an advantage over vanilla DPO, which can overfit to a narrower distribution. Meanwhile, $\beta$-DPO [141] adaptively tunes the regularization term based on the data quality, making it more robust to noisy preferences. Stepwise DPO (sDPO) [142] partitions the preference dataset to perform incremental updates, leveraging a stronger intermediate reference model at each phase.
DPO methods are particularly advantageous for structured problem-solving, like in creative writing or complex reasoning because they can directly incorporate human preferences and avoid undesired behavior without heavily relying on large-scale reward modeling or complex RL training loops [139]. However, a recurring drawback is their sensitivity to distribution shifts, e.g., when the model starts generating out-of-domain responses, alignment performance can drop unless the reference model or preference data is iteratively updated [143]. Moreover, purely relying on pairwise or setwise human judgments can still introduce label noise or ambiguity, especially for creative or unstructured tasks [144]. Despite these limitations, DPO-based techniques are promising for balancing helpfulness and correctness in open-ended LLM outputs  [145].
@@
Actor-critic methods, such as Advantage Actor-Critic (A2C) and Deep Deterministic Policy Gradient (DDPG), have been effectively adapted to optimize prompts for LLMs. Frameworks like Prompt Actor-Critic Editing (PACE) [151] employ an iterative process where the actor (the LLM) generates a response $a$ based on a prompt $p$ and input $X$. This process is formalized as
@@
    Adaptive process: It includes the sequence of automated activities performed to create an implementation. It requires devising new means to predict the expected outcomes of the available activities, selecting and adapting an activity to the current context, understanding the degree to which the sequence advances towards creating an implementation, and learning new knowledge available during the process. Also, when addressing collaboration between humans and LLMs to tackle unexpected challenges, such as handling zero-day attacks, the process necessitates reasoning, understanding of prior instructions, and intuitive decision-making within the context of new parameters and constraints. To automate this process, exploring reasoning techniques, including deductive reasoning, inductive reasoning, analogical reasoning, common sense reasoning, tree-of-thoughts, multiple chains of thought, causal reasoning, heuristic reasoning, and symbolic reasoning, is required. Among these, the primary human thought process often involves mapping the current problem to a previously encountered one or identifying similarities with analogous problems, like in analogical reasoning. Consequently, an effective approach to problem modeling could involve neuro-symbolic representations that allow LLMs to dynamically learn and adapt in real-time. Techniques such as grokking [164], which enable models to discover relationships and patterns through iterative refinement, and masked LLMs are promising methods to achieve this goal. These approaches empower the model to derive connections on the fly, effectively merging learned representations with reasoning capabilities.